OpenStudy (anonymous):

how do you graph x=e^-t and y=e^3t

OpenStudy (campbell_st):

its a set of parametric equations so take the 1st equation and make t the subject. then substitute into the 2nd equation... and you'll have something you can plot

OpenStudy (anonymous):

i got y=x^-3 but the values i got on the calculator didn't look like the graph that my teacher had posted

OpenStudy (amistre64):

x=e^-t and y=e^3t hmm, are we able to convert the first into the second ..... x = e^-t x e^(4t) = e^-t e^(4t) x e^(4t) = e^(3t) lol, that didnt work out like i wanted it too